# On Future Power System Digital Twins: A Vision Towards a Standard Architecture

### v2

@@ I Introduction
This paper begins by examining the evolution of the DT concept across various engineering domains before narrowing its focus to the power systems domain. Then, it identifies and highlights key features that we consider fundamental for the future development of DTs in power systems. By reviewing different definitions and capabilities, we propose a new standard definition of DTs specifically tailored to power systems. Building on the proposed definition, a vision of a standard DT ecosystem architecture designed to offer services beyond real-time monitoring, control, and operation is proposed. Later, we elaborate on our vision for DTs in advanced power system life cycle phases (i.e., long-term planning), highlighting challenges that remain, such as managing measurement and model errors, and uncertainty propagation. We also present our vision of how artificial intelligence (AI) and machine learning (ML) can enhance several power systems DT components and modules established in the proposed standard ecosystem architecture. Finally, we outline a structured approach for implementing the proposed DT ecosystem architecture. In summary, the main contributions of this paper are as follows:
@@
  • Addressing power systems DTs as a large-scale software development challenge: Our proposed envisioned architecture addresses the large-scale software development challenges often overlooked by existing DT platforms, advocating for a standardized approach to manage the complexity and scale of modern power systems. The proposed power system DT architecture integrates advanced data management capabilities to address measurement and model errors and uncertainties, enhancing the accuracy and reliability of the DT for robust decision-making support.
  • Power systems DTs as a System-of-Systems approach: We adopt a system-of-systems perspective, allowing for integrating DTs developed at various levels, ensuring scalability and flexibility in more complex systems.
@@ IV-A Life Cycle Aligned With TSOs’ & DSOs’ Processes
Fig. 5 attempts to reconcile the time dimension of power system DTs with the DSOs needs of integrating multiple models used for different operation and planning tasks. By highlighting data (and information) flows between models used for tasks located at different time horizons (e.g., seconds, minutes, years, decades), the envisioned DT architecture and its key components are linked to the continuous cycle of processes performed by the system operators. Nevertheless, integrating data with different time resolutions within these processes and their underlying simulation models poses significant challenges. For example, consider the Dutch household’s demand profile with several time resolutions presented in Fig. 6. Using demand profiles with different time resolutions as input for the DT simulation models can alter their output due to the inherent information loss, especially considering the stochastic behavior loss in lower time resolution data. As a result, important dynamics, such as peak demands, can be overlooked, impacting the power system’s operational resilience, for instance, by underestimating the peak consumption. Therefore, selecting the appropriate time resolution for each simulation model, performing data downsampling or upsampling as needed, and ensuring time synchronization of measurement data and simulation models are essential for accurate results. Nevertheless, this research area is usually overlooked in the power system community.
@@ IV-B Life Cycle Challenges
IV-C Expected Benefits for TSOs & DSOs
We propose that the envisioned DT standard definition and ecosystem architecture offer multifaceted benefits for TSOs and DSOs. The advanced data management backbone ensures robust data engineering and analytics capabilities within a single platform, providing TSOs and DSOs with enhanced insights into power system control and operation. Integrating high-fidelity simulation models allows system operators to effectively manage error and uncertainty propagation, facilitating comprehensive analyses across all control, operation, and planning components. This setup enables the development and validation of new services, which can be seamlessly redeployed at the system operator control center premises. Integrating a dedicated communication component enhances coordination and safeguards platform integrity, crucial for mitigating cyberattack risks when interfacing with external parties. Adopting multi-DT systems supports long-term infrastructure planning for combined power systems. By consolidating all simulation models, data analytics, data engineering, operation and planning components, and Grid as a Service into a single ecosystem, system operators can efficiently test, validate, and implement additional modules with minimal effort. Once the backbone is established, managing and expanding the system can become a straightforward procedure.
@@ V The Role of Artificial Intelligence & Machine Learning in Future Power System Digital Twins
VI Steps to Implementation
A structured approach is necessary to implement the envisioned DT ecosystem architecture effectively. First, establish a collaborative framework by forming a consortium that includes TSOs, DSOs, software engineers, computer scientists, and academic researchers. The consortium’s goal should be to develop mature system engineering design methodologies for DT implementation. Second, standardize workflows and data sharing by creating data protocols and APIs to facilitate communication between DT modules and physical systems. Ensure that all relevant parties consistently use data formats like JSON, XML, and CSV, and that asset information adheres to industry standards such as CIM. Third, the data engineering and analytics components must be developed and implemented. This includes data formatting, cleaning, integration, storage, and visualization capabilities. Utilize advanced AI and ML models to enhance data processing, visualization, anomaly detection, and data imputation. Once the system’s backbone is developed, system operators can begin implementing their operation and planning modules. The next step is validation and testing, where the DT ecosystem is tested in a controlled environment using historical data and simulated scenarios to ensure proper operation.
VII Conclusion

### v3

@@ I Introduction
This paper begins by examining the evolution of the DT concept across various engineering domains before narrowing its focus to the power systems domain. Then, it identifies and highlights key features that we consider fundamental for the future development of DTs in power systems. By reviewing different definitions and capabilities, we propose a new standard definition of DTs specifically tailored to power systems. Building on the proposed definition, a vision of a standard DT ecosystem architecture designed to offer services beyond real-time monitoring, control, and operation is proposed. Later, we elaborate on our vision for DTs in advanced power system life cycle phases (i.e., long-term planning), highlighting challenges that remain, such as managing measurement and model errors, and uncertainty propagation. Finally, we present our vision of how artificial intelligence (AI) and machine learning (ML) can enhance several power systems DT components and modules established in the proposed standard ecosystem architecture. In summary, the main contributions of this paper are as follows:
@@
  • Addressing power systems DTs as a large-scale software development challenge: Our proposed envisioned architecture addresses the large-scale software development challenges often overlooked by existing DT platforms, advocating for a standardized approach to manage the complexity and scale of modern power systems. The proposed power system DT ecosystem architecture integrates advanced data management capabilities to address measurement and model errors and uncertainties, enhancing the accuracy and reliability of the DT for robust decision-making support. Moreover, we adopt a system-of-systems perspective, allowing for integrating DTs developed at various levels, ensuring scalability and flexibility in more complex systems.
@@ IV-A Life Cycle Aligned With TSOs’ & DSOs’ Processes
Fig. 5 attempts to reconcile the time dimension of power system DTs with the DSOs needs of integrating multiple models used for different operation and planning tasks. By highlighting data (and information) flows between models used for tasks located at different time horizons (e.g., seconds, minutes, years, decades), the envisioned DT architecture and its key components are linked to the continuous cycle of processes performed by the system operators. Nevertheless, integrating data with different time resolutions within these processes and their underlying simulation models poses significant challenges. For example, consider the Dutch household’s demand profile with several time resolutions presented in Fig. 6. Using demand profiles with different time resolutions as input for the DT simulation models can alter their output due to the inherent information loss, especially considering the stochastic behavior loss in lower time resolution data. As a result, important dynamics, such as peak demands, can be overlooked, impacting the power system’s operational resilience, for instance, by underestimating the peak consumption. Therefore, selecting the appropriate time resolution for each simulation model, performing data downsampling or upsampling as needed, and ensuring time synchronization of measurement data and simulation models are essential for accurate results. A multi-domain novel generation modelling (NGM) platform proposed in [54] supports multi-time-scale operation within the same system. Nevertheless, this research area is usually overlooked in the power system community.
@@ V The Role of Artificial Intelligence & Machine Learning in Future Power System Digital Twins
VI Conclusion